59. f(x) = 5 + 38 = UL UNIL 1,1] 60. f(x) = 2x6 – 15x4 + 24x² on (-2, 27 61 fly) = -nn -2.21 QUESTION 60.
determine the location and value of the absolute extreme values of f on the given interval, if they exsist.
Answer #1

og flm) = 226-15*4 +2422 on [-2,2] critical point = f(n): 1225–60x +48x = 0 (65_583+4x)=0 X(a4-5**+ 4) = 0 x=0, (0-1)(3+4)=0